A father is four times as old as his son and his daughter is five years younger than her brother. If their combined age is 73 years, find the age of each.
Answer by rfer(16322)

x=son
4x=father
x-5=daughter
--------------
x+4x+x-5=73
6x=78
x=78/6
x=13
4x=52
x-5=8